OpenStudy (anonymous):

tan2k=tank what is k?

OpenStudy (anonymous):

k=0

OpenStudy (anonymous):

npi is better ?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

@DHASHNI I wanna see ur solution,pls. (:

OpenStudy (anonymous):

tan(2x) - tan x = 0 2tan x / (1 - tan^2(x)) - tan x = 0 tan x [ 2 / (1 - tan^2(x)) - 1 ] = 0 tan x [ 2 - 1 + tan^2(x) ] = 0 tan x (1 + tan^2(x)) = 0 tan x = 0 x = 0, π